How many litres of a 68% concentration nitric acid should be diluted with 400 litres of water to give a solution with 3% concentration of nitric acid??

Answer by richwmiller(17219)

from view of water
.32x+400=.97(400+x)
.32x+400=388+.97x
12=.65x
x=18.461538461538461538461538461538 litres of 68% concentration nitric acid
check
5.9076923076923076923076923076923+400=388+17.907692307692307692307692307692
405.90769230769230769230769230769=405.90769230769230769230769230769
ok
view of acid
.68x=.03(x+400)
.68x=.03x+12
.65x=12
x=18.461538461538461538461538461538 litres of 68% concentration nitric acid